# Writing – original draft, Data curation, Formal analysis, Methodology, Conceptualization

### v1

@@ 2.9.1 Disease progression in vectors
As mentioned in Section 2.4, we do not consider individual vectors separately. Instead, we used a patch to represent vector density and employed a stochastic process to model disease progression. Specifically, we focused on three disease states of the vectors: susceptible, exposed, and infected, as illustrated in Fig. 14.
@@
Each patch $k$ is initialized with counts for susceptible ($S_{v}^{k}$), exposed ($E_{v}^{k}$), and infected ($I_{v}^{k}$) vectors, ensuring that these values remain below a defined vector carrying capacity specific to each patch. The temporal dynamics of vector density within a patch are modeled using a system of differential equations. The changes in vector counts for a given patch $k$ are determined as follows. Factors with superscript $k$ are unique to a given patch, while others are patch-independent.
$\frac{dS_{v}^{k}}{dt}=h_{v}^{k}-\lambda_{v}^{k}S_{v}^{k}-\mu_{v}S_{v}^{k}$ (9)
$\frac{dE_{v}^{k}}{dt}=\lambda_{v}^{k}S_{v}^{k}-v_{v}^{k}E_{v}^{k}-\mu_{v}E_{v}^{k}$ (10)
$\frac{dI_{v}^{k}}{dt}=v_{v}^{k}E_{v}^{k}-\mu_{v}I_{v}^{k}$ (11)
where:
  • $h_{v}^{k}$: Total birth rate of vectors
  • $\lambda_{v}^{k}$: Average infection rate of vectors
  • $v_{v}^{k}$: Average rate of progression of vectors from exposed state to infectious state
  • $\mu_{v}$: Average death rate of vectors
Figure 14: Mosquito state transition.
The average infection rate of vectors in patch $k$, $\lambda_{v}^{k}$, is calculated as follows:
@@
where:
  • $b_{v}^{k}$: The number of bites per vector per unit time
  • $\beta_{vh}$: The probability of transmission from an infectious agent to a susceptible vector given that contact between the two occurs
  • $I_{h}^{k}$: The number of infected agents in the patch $k$
  • $N_{h}^{k}$: Total number of agents in patch $k$
The number of bites per vector per unit of time in patch $k$ which is defined in the following way:
@@
where:
  • $b^{k}$: The total number of contacts between agents and vectors (bites) in patch $k$ per unit time
The total number of contacts between agents and vectors (bites) in patch $k$ is defined in the following way:
@@
where:
  • $\sigma_{v}$: The maximum number of bites per vector per unit time.
  • $\sigma_{h}$: The number of bites a human can get per unit of time.
@@
where:
  • $tinc_{v}^{k}$: The incubation time of the virus in the mosquito in patch $k$. This depends on the temperature of the patch ($T^{k}$) and is calculated as follows:
    • If $T^{k}<21$, then $tinc_{v}^{k}\sim U(10,25)$
    • If $2621\leq T^{k}<$ 26, then $tinc_{v}^{k}\sim U(7,10)$
    • If $3126\leq T^{k}<$ 31, then $tinc_{v}^{k}\sim U(4,7)$
The parameters used for dengue transmission [38] are provided in Table. 4, with rate values specified on a per-day basis. Since the model includes multiple vector patches, a range of values was selected for some parameters, indicating minimum and maximum values (enclosed in brackets).
Table 4: Parameters for dengue simulation
Parameter Value/Range
$\psi_{v}$: Average natural emergence rate of mosquitoes 0.3
$\mu_{v}$: Average death rate of vectors 1/14
$\sigma_{v}$: The maximum number of bites per vector per day 0.5
$\sigma_{h}$: The number of bites a human can get per day 10
$K_{v}$: Mosquito carrying capacity [100-200]
$\beta_{vh}$: The probability of transmission from an infectious agent to a susceptible mosquito 0.33
$\beta_{hv}$: The probability of transmission from an infectious mosquito to a susceptible agent 0.33
2.9.2 Disease progression in agents
The infection rate of agents from infectious mosquitoes is determined by the frequency of bites an agent receives, the likelihood that the bite is from an infectious mosquito, and the probability that the infection will be successful given the bite. This rate can be represented as:
@@ 2.9.2 Disease progression in agents
where:
  • $b_{k,h}$: the number of bites an agent receives per unit time from mosquitoes in patch $k$
  • $\beta_{h}$: the probability of successful infection given a bite from an infectious mosquito
  • $I_{k}^{v}$: the number of infectious mosquitoes in patch $k$
  • $N_{k}^{v}$: the total number of mosquitoes in patch $k$
To calculate the probability of a susceptible agent becoming infected, the infection rate must be converted into a probability of infection within a given time step. We run the simulation and update agent movements in time steps of $\Delta t$, while the disease states of agents and vectors are updated every $n\cdot\Delta t$ time steps. In our simulation, agent movements are updated every 1 minute, and disease statuses are updated every 5 minutes. Assuming the time to infection follows an exponential distribution [14], the probability of infection at the end of the time interval $\Delta t$ is given by:
@@
where:
  • $p_{k}$: infection probability of a susceptible person in patch $k$
  • $\lambda_{k,h}(t)$: infection rate of an agent
Then, a random variable between 0 and 1 is generated every 5 minutes to check if it falls above the infection probability. If the random variable exceeds the probability, the agent is marked as exposed. Once an agent is exposed, the disease transitions follow the Markov model as shown in Fig. 13.
An infected person may either stay asymptomatic, develop mild symptoms, or, in rare cases, progress to severe symptoms [38]. Many people with dengue remain asymptomatic, showing no visible symptoms. Those with mild symptoms experience typical dengue signs like fever, headache, and body aches, which often resolve with basic supportive care, allowing for full recovery. Severe dengue, however, can bring serious symptoms like intense abdominal pain, persistent vomiting, bleeding, or shock, often requiring hospitalization and posing life-threatening risks.
The likelihood of transitioning between these states depends on various factors, including age, prior dengue infections, and the quality of medical care. For instance, the likelihood of an infected person developing symptoms may vary based on their immune response.
The Markov model outlined in Section 2.7 is used to determine state transitions. For dengue, the model focuses only on mild and severe symptomatic states. By combining probabilistic state transitions with random transition periods, the model can effectively simulate diverse disease progression outcomes among individuals.
2.9.3 Hospitalization mechanism
There is a high probability of symptomatic patients visiting the hospital. At the start of each day in the simulation, all agents’ disease statuses are checked. If symptomatic agents are present, a random hospitalization probability is generated. If this probability exceeds a certain threshold, the agent is hospitalized. For severe cases, this threshold is lower, reflecting the higher likelihood of hospitalization. In our simulation, we used 0.4 for mild cases and 0.9 for severe cases.
Once the agents requiring hospitalization are identified, they are removed from both the environment and the patches and placed in a virtual space. Hospitalized agents do not contribute to further disease spread. However, their disease states continue to update as outlined in Fig. 13. Eventually, these agents either recover or die. Recovered hospitalized agents are returned to both the environment and their respective patches.
2.9.4 Intervention mechanism
@@
3 Results and discussion
3.1 Spectral clustering results
Distinct behavioral patterns within each occupation class were derived by applying the Spectral Clustering algorithm to each occupation class. This was done by manual inspection of the output of the clustering algorithm. Table  5 presents the synthesized results. Note that normal behavior refers to a typical workday for each class.
Table 5: Distinct behavioral patterns Obtained for occupation classes through spectral clustering method.
@@ 3.1 Spectral clustering results
Animal Farmer Work area is located in close proximity to the residence.
Work area is situated separately from the residence.
Individual resides and works within a designated Agricultural Zone.
Bank Worker Normal workday behaviour
Weekend and holiday behaviour
@@
Night shift at the hospital with daytime duties at a school medical center
Farmer Work area is located in close proximity to the residence.
@@
Garment Admin Weekend and holiday behaviour
Normal workday behaviour
@@
Sales Rep Normal workday behaviour
Holiday behaviour
@@
Works from morning to afternoon & evening to night with a break in between. also includes holiday behavior
@@
Normal teacher behaviour (travelling from home)
@@
After examining the results, it can be observed that the algorithm excels at producing clusters with distinct behavioral patterns, ranging from regular work routines to holiday behaviors in alignment with real-world intuition. Furthermore, the algorithm effectively filters out anomalous patterns, as demonstrated in the instances above. The final results and the human motion dataset we created to fuel the AVSim can be found online at [44].
3.2 Validation of human motion modeling
@@
To validate the accuracy of the human motion model, 10000 samples of daily person trajectories were generated for each profession class. Then, the probability distributions were reconstructed using them to compare with the original probability distributions. These synthetic and original plots can be seen in Fig. 15. We can see that the synthetic distributions closely resemble the ground truth distributions, thus proving the above method to be highly accurate in replicating human behavior.
Figure 15: Validation of bimodality.
3.3 Air borne disease simulation results
In this sector, a few simulations were conducted to research air-borne disease spread. However, it is important to note that these results are only aligned through a subset of possible research directions. One can perform various analyses based on their need and interests. The simulations were conducted in a simulated environment replicating the Kandy district.
3.3.1 Contact tracing
By default, the AVSim will generate detailed descriptions of each agent in the simulation, such as daily routines, daily person trajectories, disease progression history, and daily contacts. This is illustrated in Fig. 16. The AVSim can be customized to generate any other required information.
Figure 16: Contact tracing capability of AVSim.
@@ 3.3.1 Contact tracing
Figure 17: All the contacts in an uncontrolled, i.e., no vaccine nor quarantine disease-transmitting moment for 50 days simulation in a part of "KandyCity" and "Pallekele." Initially, three students were infected randomly. The receiving agent is shown by filled circles and the cause; the transmitting agent is shown by outer borders in each instance.
Figure 18: Agents count in different states in an uncontrolled epidemic outbreak, which started with three infected students.
The ability to trace back contacts is a key feature that is crucial in any type of ABM. Fig. 17 shows the locations where the disease was transmitted from an infected agent to a healthy agent. This is essentially the contact tracing mechanism of the ABM in action. When analyzing, it can be seen more contacts are present in the "Residential Zones." This can be described as the area of houses compared to other places being lower. Therefore, there is a high possibility of more contact in houses. Also, another major observation is that although there are higher contacts in residential zones, the spread of disease to other communities, i.e., residences present in other cities, is by contacts in work locations or outside. This can be verified by Fig. 16. Starting with one infected student agent, the disease transmits to another student agent at the school. Then, the received student agent transmits the disease to the nurse at the student agent’s home. This leads to the spread of disease in a hospital, which is the working location of the nurse. Likewise, it can be observed that there are several branches of contact between the second student who was contacted and the initial student at the school. It proves work locations (here, a school for the student and a hospital for the nurse) are involved in spreading diseases to different communities through the agents in the same work locations. Therefore, identifying these kinds of contact patterns and predictions at early stages will enable policymakers to pinpoint high-risk areas and implement targeted interventions to prevent the spread of the disease in these regions.
Figure 19: Disease spread in different professional classes in an uncontrolled epidemic outbreak which started with three infected students.
@@
Figure 20: Disease spread in different professional classes in an uncontrolled epidemic outbreak which started with three random infected agents.
3.3.2 Disease progression in an uncontrolled environment
@@
This AVSim simulation can be executed without imposing policies to reduce the spread, which is known as an uncontrolled environment. Fig.  18 shows an uncontrolled epidemic outbreak where three random student agents were initially infected. We can see that this plot closely resembles the theoretical SEIR model curve, hence validating the practical approach of AVSim.
3.3.3 Effect of initially infected agents on disease progression
The AVSim allows the analysis of disease spread patterns when agents from a specific class were initially infected. To demonstrate this, three students were initially infected in an uncontrolled environment, and then the disease spread to other classes over the simulation days was observed. The Observations are shown in Fig.19, highlighting only the top three classes that had the highest percentage of infected agents. For approximately up to 50 days, a considerable percentage of the disease was spread to teachers, w.r.t. other classes. This is reasonable because contact from students is mostly present in schools, which is the work location of teachers. This analysis helps identify the most vulnerable groups to the disease concerning initial infection.
Furthermore, it can be observed that the spread to other classes in Fig. 19 becomes higher when the days pass, showing the effect of contacts in an uncontrolled environment. When the agents are randomly infected, it is hard to predict a pattern of spreading to other classes, and it depends on contact tracing. For example, in the simulation provided by Fig  20, when the agents were randomly selected, initially infected agents were one student, one doctor, and one teacher. However, in that case, disease spread had a significant impact on the medical sector agents. One key observation from the two simulations mentioned above is that the student class is consistently more affected. This is because most agents in different professional classes have student agents in their residences, where they will interact more with student agents. Therefore, more controlled strategies should be implemented to minimize the risk of disease spread to other classes from these vulnerable groups.
3.3.4 Quarantine and vaccination measures
Figure 21: Effect of quarantine mechanism in airborne diseases w.r.t. uncontrolled case.
Figure 22: Effect of vaccine mechanism in airborne diseases w.r.t. uncontrolled case.
Figure 23: Effect of quarantine and vaccination mechanism in airborne diseases.
@@ 3.3.4 Quarantine and vaccination measures
We have tested several controlled mechanisms to reduce the spread of airborne diseases. The following conclusions were assumed based on the parameters that we set as described in the 2.8 section.
  1. A
    Effect of Quarantine: Fig. 21 shows the number of infected agents in different quarantine mechanisms compared to uncontrolled cases. The AVSim allows two mechanisms in quarantine: quarantine by profession class and self-quarantine of individuals when PCR results are positive. It can be observed that when combined mechanisms are used, the peak of infected agents is reduced by approximately 87.5%.
  2. B
@@
    Effect of Vaccine: Fig. 22 shows the number of infected agents in different vaccine mechanisms compared to uncontrolled cases. We analyzed the effect of vaccinating all the agents when the infected percentage of the population is greater than 5% and 10%. It can be deduced that early vaccination events cause a higher disease spread reduction.
  3. C
    Controlled vs. Uncontrolled cases: Fig. 23 shows the number of infected agents in controlled via vaccination or quarantine and uncontrolled cases. It can be observed that a higher reduction in disease spread was obtained from proper quarantine mechanisms.
3.4 Vector-borne disease simulation results
The simulations were conducted with 2000 individual agents, using Kandy District as the simulation environment. The transportation model was not included in the simulation, based on the assumption that vector bites do not occur in public transportation.

### v2

@@ III-H1 Disease progression in vectors
As mentioned in Section III-E, we do not consider individual vectors separately. Instead, we used a patch to represent vector density and employed a stochastic process to model disease progression. Specifically, we focused on three sequential disease state transitions of the vectors: susceptible, exposed, and infected.
@@
Each patch $k$ is initialized with counts for susceptible ($S_{v}^{k}$), exposed ($E_{v}^{k}$), and infected ($I_{v}^{k}$) vectors, ensuring that these values remain below a defined vector carrying capacity specific to each patch. The temporal dynamics of vector density within a patch $k$ are modeled using a system of differential equations as follows.
$\frac{dS_{v}^{k}}{dt}=h_{v}^{k}-\lambda_{v}^{k}S_{v}^{k}-\mu_{v}^{k}S_{v}^{k}$ (8)
$\frac{dE_{v}^{k}}{dt}=\lambda_{v}^{k}S_{v}^{k}-v_{v}^{k}E_{v}^{k}-\mu_{v}^{k}E_{v}^{k}$ (9)
$\frac{dI_{v}^{k}}{dt}=v_{v}^{k}E_{v}^{k}-\mu_{v}^{k}I_{v}^{k}$ (10)
Here, $h_{v}^{k}$ is the total vector birth rate, $\lambda_{v}^{k}$ is the average infection rate, $v_{v}^{k}$ is the rate of progression from exposed to infectious, and $\mu_{v}$ is the average vector death rate. The $\lambda_{v}^{k}$, is calculated as follows:
@@
where $b_{v}^{k}$ is the number of bites per vector per unit time, $\beta_{vh}$ is the probability of transmission from an infectious agent to a susceptible vector upon contact, $I_{h}^{k}$ is the number of infected agents in patch $k$, and $N_{h}^{k}$ is the total number of agents in that patch. The $b_{v}^{k}$, is calculated as:
@@
where $b^{k}$ is the total number of vector-human contacts (bites) per unit time in patch $k$.
@@
Here, $\sigma_{v}$ is the maximum number of bites per vector per unit time, $\sigma_{h}$ is the number of bites a human can receive per unit time, $N_{v}^{k}$ is the number of vectors, and $N_{h}^{k}$ is the number of agents in patch $k$.
@@
where $tinc_{v}^{k}$ is the temperature-dependent incubation time of the virus in mosquitoes within patch $k$. Based on the patch temperature $T^{k}$, $tinc_{v}^{k}$ is drawn from a uniform distribution: $U(10,25)$ if $T^{k}<21$, $U(7,10)$ if $21\leq T^{k}<26$, and $U(4,7)$ if $26\leq T^{k}<31$.
The parameters used for dengue transmission were taken from similar studies [8] and the WHO website [35]. Since the model includes multiple vector patches, a range of values was selected for some parameters to add variability.
III-H2 Disease progression in agents
The infection rate of agents, $\lambda_{k,h}(t)$ in patch $k$ depends on the number of bites per unit time $b^{k}_{h}$, the proportion of infectious mosquitoes, ${I_{k}^{v}}/{N_{k}^{v}}$, and the probability of successful transmission, $\beta_{hv}$. This relationship is given by:
@@ III-H2 Disease progression in agents
Probability of an exposed agent becoming infected within a given time step is calculated as follows. The simulation is executed in time steps of $\Delta t$ (1 min in our case), while the disease states of agents and vectors are updated every $n\cdot\Delta t$ (5 min) time steps. Assuming the time to infection follows an exponential distribution [8], the probability of infection at the end of the time interval $\Delta t$ is given by:
@@
Once an agent is exposed, the disease transitions follow the Markov model as shown in Fig. 8. An infected person may either stay asymptomatic, develop mild symptoms, or, in rare cases, progress to severe symptoms [35]. The likelihood of transitioning between these states depends on various factors, including age, prior dengue infections, and the quality of medical care. For instance, the likelihood of an infected person developing symptoms may vary based on their immune response.
The Markov model outlined in Section III-F is used to determine state transitions. For dengue, the model focuses only on mild and severe symptomatic states. By combining probabilistic state transitions with random transition periods, the model can effectively simulate diverse disease progression outcomes among individuals.
III-H3 Hospitalization mechanism
There is a high probability of symptomatic patients visiting the hospital. At the start of each day in the simulation, all agents’ disease statuses are checked. If symptomatic agents are present, they will be hospitalized based on some hospitalization probability. For severe cases, this probability is higher.
Once the agents requiring hospitalization are identified, they are isolated from other agents and put into hospitals. It is assumed that hospitalized agents do not contribute to further disease spread. However, their disease states continue to update as outlined in Fig. 8. Eventually, these agents either recover or die. Recovered hospitalized agents are returned to both the environment and their respective patches.
III-H4 Intervention mechanism
@@
IV Results and discussion
IV-A Spectral clustering results
As discussed in  III-A2, distinct behavioral patterns within each occupation class were derived by applying the Spectral Clustering algorithm to each occupation class. This was done by manual inspection of the output of the clustering algorithm. Table II presents the synthesized results. Note that normal behavior refers to a typical workday for each class.
TABLE II: Distinct behavioral patterns obtained for selected occupation classes through the spectral clustering method.
@@ IV-A Spectral clustering results
Night shift at hospital, daytime at school medical center
Farmer Work area is near the residence
@@
Works morning to afternoon and evening to night, with a break in between; includes holiday behavior
@@
Normal behaviour of a teacher travelling from home
@@
After examining the results, it can be observed that the algorithm excels at producing clusters with distinct behavioral patterns, ranging from regular work routines to holiday behaviors in alignment with real-world intuition. The final results and the human motion dataset we created to fuel the AVSim can be found online at [38].
IV-B Validation of human motion modeling
@@
To validate the accuracy of the human motion model, 10000 samples of daily person trajectories were generated for each profession class. From this synthetic data, the probability distributions for showing up at different locations with time were reconstructed to compare with the original ground truth probability distributions. These synthetic and original plots for the doctor and teacher classes can be seen in Fig. 9. We can see that the synthetic distributions closely resemble the ground truth distributions, thus proving the above method to be highly accurate in replicating human behavior.
Figure 9: Validation of our motion model (Synthetic) against real world motion data (Ground truth).
IV-C Airborne disease simulation results
@@
Figure 10: All the contacts in an uncontrolled simulation for 50 days in a part of ”Kandy City” and ”Pallekele”. Initially, three students were infected randomly. The receiving agents are shown by filled circles and the transmitting agent is shown by outer borders in each instance.
In this section, we discuss the simulation results to research air-borne disease spread. However, it is important to note that these results only constitute a subset of possible research directions. One can perform various analyses based on their need and interests. All these simulations consisted of approximately 2000 agents. This is just for demonstration and, based on the magnitude of the simulation environment, the agent count could be increased at the cost of computational power.
IV-C1 Tracing path of disease spread
@@
Figure 11: (a) Agents count in different states in an uncontrolled epidemic outbreak which started with three infected students, Disease spread in different professional classes in an uncontrolled epidemic outbreak which started with (b) three infected students, (c) three random infected agents, (d) Effect of vaccine mechanism in airborne diseases w.r.t. uncontrolled case, (e) Effect of quarantine and vaccination mechanism in airborne diseases, (f) Effect of quarantine and vaccination mechanism in airborne diseases.
A key ABM feature is contact tracing, shown in Fig. 10, which highlights disease transmission locations. Residential zones show more contacts, likely due to higher agent density. However, disease spread across communities often occurs through work locations. For example, in Fig 12 an infected student transmits the disease at school, which then reaches a nurse at home, leading to further spread at the hospital where the nurse works. This illustrates how workplaces link communities in transmission chains. Leveraging AVSim, the early identification of such patterns enables policymakers to target high-risk zones and apply timely interventions.
AVSim also generates detailed descriptions of each agent, including daily routines, person trajectories, disease progression histories, and daily contacts. This is illustrated in Fig. 12.
Figure 12: Tracing the paths of contacts of AVSim.
@@
AVSim simulation system can be executed in an uncontrolled environment without imposing any policies that reduce the spread. Fig. 11a shows such an epidemic outbreak where three random student agents were initially infected. We can see that this plot closely resembles the theoretical SEIR model curve [39], hence validating the practical approach of AVSim. This can be used as the controlled experiment for the upcoming experiments.
IV-C2 Effect of initially infected agents on disease progression
AVSim enables analysis of disease spread from specific agent classes. In an uncontrolled setting where three students were initially infected, the disease was spread mostly to teachers within the first 25 days (Fig. 11b). This is due to frequent school contacts during the first couple of weeks. Over time, infections rose in other classes, revealing vulnerable groups and the effects of uncontrolled transmission. When agents are randomly infected, it is hard to predict a pattern of spreading to other classes, and it depends on the initially infected agents. For example, in the simulation provided by Fig. 11c, initially infected agents were one student, one doctor, and one teacher. For this case, the disease spread had a significant impact on the medical sector agents. However, students were consistently among the most affected in all simulations, likely due to their presence in many households.
This demonstrates that AVSim can identify which professional classes are most vulnerable based on the outbreak’s initial state, enabling the design of targeted control strategies to limit disease spread from these high-risk groups.
IV-C3 Study of Quarantine and vaccination measures
@@
We have tested several controlled mechanisms to reduce the spread of airborne diseases. The following conclusions were assumed based on the parameters that we set as described in the Section III-G.
  1. 1.
@@
    Effect of Vaccine: Fig. 11d shows infections under different vaccine strategies versus no control. We analyzed the effect by vaccinating all the agents when the infected percentage of the population is greater than 5% and 10%. It can be deduced that early vaccination leads to greater spread reduction [40].
  2. 2.
    Effect of Quarantine: Fig. 11e compares infected agent counts under different quarantine mechanisms versus no control. AVSim supports two types: profession-based quarantine and self-quarantine upon positive PCR. We find that when combined, they reduce the infection peak by  87.5
  3. 3.
    Vaccination and Quarantine hybrid: Fig. 11f compares infections when vaccination and quarantine were applied independently and together. Quarantine alone reduced infections more, likely due to complete movement restrictions in the simulation. However, this idealized model doesn’t reflect real-world conditions, where quarantined individuals still require some contact for basic needs, potentially leading to higher actual transmission risk.
The results also show that vaccination helps control the spread more quickly. An ideal strategy would combine targeted vaccination, prioritizing high-risk groups like health professionals, with effective quarantine of infected agents and areas. Since AVSim is fully customizable, it can be adapted to simulate more realistic quarantine and vaccination scenarios tailored to specific geographic or contextual settings, providing policymakers with more applicable insights.
IV-D Vector-borne disease simulation results
Following similar logic to section  IV-C, all simulations were conducted with 2000 individual agents.